# Theory uncertainties of the irreducible background to VBF Higgs production

### v1

@@ III.1 Observable definitions
Additional data are collected on https://doi.org/10.5281/zenodo.17015614, together with the Rivet Buckley et al. (2013); Bierlich et al. (2020, 2024) analysis routine. The Higgs boson is required to be within the central region of the detectors () in association with at least two accompanying jets. The jets are clustered using the FastJet implementation of the anti- algorithm Cacciari et al. (2012), and must satisfy the following cuts:
@@
We primarily consider jets with a radius of 0.4, but also provide results for radii of 0.3, 0.6 and 0.8, to quantify the impact on selected differential observables.222Our study of the VBF Higgs boson production process in Ref. Buckley et al. (2021) examined jet -values from 0.3 to 1.0, and indicated a strong -dependence due to the kinematic requirements on the two jets. The R-dependence increased upon the imposition of the VBF phase space cuts, so it is important to check if the same effects occur for the background. Note that the two leading jets in VBF Higgs boson production are primarily quark jets, while in ggF they are primarily gluon jets. This may led to some differences in the -dependence. Jets are ranked by the magnitude of their transverse momentum.
@@
To distinguish between normalization and shape differences of the predictions, all results will be normalized to the inclusive cross section for the final state, defined at the particle level. This means in particular that a normalization difference can occur through the different behavior of the parton shower simulation, for example if the parton shower shifts a jet above/below the cut by means of additional radiation and/or recoil effects. The cross section is inferred from a relatively inclusive distribution, which we choose to be the Higgs boson rapidity distribution in the two jet final state.
III.2 Results of the comparison
@@
We are now ready to compare results from the MiNNLO simulation as used by ATLAS and CMS to those from the tools introduced in Sec. II. The MiNNLO calculation is matched to the default Pythia parton shower (“simple shower”), both with and without the “dipole-recoil” option, and to the Herwig angular-ordered parton shower. The results are compared to a MEPS@NLO simulation from Sherpa, including the , , and final states at NLO precision, and up to at LO accuracy. This sample will be dubbed MEPS@NLO in the following. All predictions are at parton level, without multi-parton interactions and non-perturbative corrections. We also use a fixed-order NNLO calculation from NNLOJet as a reference. As a reminder, MiNNLO is NNLO accurate for the inclusive Higgs boson cross section, NLO accurate for the 1-jet final state, and LO accurate for the 2-jet final state.
@@
Figure 1 shows these comparisons. The middle panels of each sub-figure display the ratio to the inclusive fixed-order calculation from NNLOJet, while the bottom panels compare the event-generator predictions, normalized to the MiNNLO + Pythia result. We find approximate agreement of the rate in the Higgs rapidity distribution of the selection (not shown here) between NNLOJet and MiNNLO + Pythia, whereas the Sherpa MEPS@NLO sample requires a scale factor of , and the MiNNLO + Herwig sample a scale factor of . In the case of the Sherpa prediction, the scale factor is consistent with the NNLO to NLO ratio of the inclusive Higgs boson production cross section Anastasiou et al. (2016), which is needed to bring the inclusive jet rates in the MEPS@NLO result into approximate agreement with the prediction from NNLOJet.
@@
This can be understood better with the help of Fig. 2, which shows the leading jet transverse momentum spectra in the inclusive and selection. Formally, the MiNNLO simulation models the first of these spectra at NLO precision, and the second at LO precision. The MEPS@NLO simulation should predict both spectra at NLO precision, but with added Sudakov suppression effects. The NNLOJet reference provides NNLO precision for the first spectrum, and NLO precision for the second. Figure 2 shows that the scaled MEPS@NLO result has the correct shape over most of the fiducial phase space, and in the medium hard region in particular. The MiNNLO predictions on the other hand agree neither with the leading jet spectrum in the selection, nor with the leading jet spectrum in the selection. In particular, we observe a significant slope difference in the medium hard region of the leading jet transverse momentum in the selection, where the MiNNLO result is formally NLO accurate. We therefore conclude that the agreement in overall rate between the MiNNLO + Pythia result and the NNLOJet prediction is rather accidental. We will discuss this point in more detail in Sec. IV.4, and in particular when comparing the distributions in Fig. 9.
@@
The rapidity separation and the invariant mass distribution of the two leading jets, shown in Figure 1, agree well between the Sherpa, Pythia, and Herwig setups. At low invariant masses, , and small rapidity separations, , all results also agree well with NNLOJet, while they predict a 20-30% increased rate in the complementary region. The distributions of the azimuthal separation generally shows good agreement between the different predictions in the region of small and intermediate separations, where they lie within a 10% envelope of each other. At large jet separations, , the MiNNLO + Pythia prediction is larger than all other predictions by about 30%, including the MiNNLO + Pythia results using the dipole-recoil option. We will discuss this in more detail in Fig. 9. The Higgs transverse-momentum distribution shows the largest shape differences between the different simulations. At low transverse momentum, , both MiNNLO + Pythia results and MiNNLO + Herwig exceed the NNLOJet result by about 20-30%, with close agreement between the MiNNLO predictions obtained with Herwig and with the Pythia dipole-recoil shower. The Sherpa MEPS@NLO prediction is about 20% lower than the NNLOJet result. Starting from , both MiNNLO + Pythia predictions agree very well with NNLOJet, whereas the Sherpa MEPS@NLO and MiNNLO + Herwig predictions create an envelope around the fixed-order result, with Sherpa MEPS@NLO being larger by about 10% and MiNNLO + Herwig being smaller by about the same amount. As can be inferred from the middle panes in all four figures, the event-generator predictions broadly lie within the seven-point scale variation uncertainty band of the fixed-order calculation, with some exceptions towards larger values of the observables.
Based on the observations in Fig. 1, one may be tempted to conclude that the modeling of the irreducible background to VBF Higgs production suffers from large theory uncertainties. This is in fact not the case. The event-generator predictions compared so far have rather different formal accuracy in the fiducial phase space, and therefore provide a different level of theoretical control over the observables of interest. Furthermore, as already discussed earlier, while in the case of NLO matching with an angular-ordered shower, truncated showers are known to yield a very small effect, an in-depth study is missing for the case in which the two hardest emissions have already been generated. Most of these effects can be mitigated once the final state is modeled consistently at NLO accuracy. This will be discussed in the next section.
@@ IV Assessment of theory uncertainties
In the following, we will use a consistent baseline for the hard, perturbative calculation – at NLO precision – and assess the influence of different theory variations on the size of the theoretical uncertainties in the fiducial phase space. Throughout the discussion, we will focus on the same four observables as in Fig. 1, with additional results available at https://doi.org/10.5281/zenodo.17015614. In Sec. IV.1, we compare parton-shower variations in NLO-matched distributions at the parton level. Sec. IV.2 focuses on matching variations in Powheg Box, while the influence of dedicated Pythia tunes and different hadronization models in Sherpa is discussed in Sec. IV.3. A study of the impact of the Higgs effective theory approximation can be found in App. A and a study of the jet-radius dependence can be found in App. B.
@@ IV.1 Parton-shower variations
As before, we scale all distributions to the NNLOJet rate based on the Higgs rapidity spectrum, yielding scale factors of in the case of Sherpa, for both Pythia showers, and for Herwig. The resulting parton-shower variations are shown in Fig. 3, including the NNLOJet NNLO prediction as a baseline. Besides the main distributions, we also show two ratio panels, with the first (middle panel) comparing the matched predictions to the NNLOJet fixed-order reference, and the second (bottom panel) showing the difference between the matched parton-shower predictions, normalized to the Powheg Box + Pythia result. All predictions in Fig. 3 acquire formal NLO accuracy in configurations.
As can be seen from Fig. 3, we find excellent, percent-level, agreement between the four matched parton-shower algorithms for the , , and distributions. For the distribution, both Pythia showers agree very well with Sherpa, while the Herwig angular-ordered shower differs from these at large transverse momentum, . In all cases, the matched parton-shower results are within the scale uncertainties of the fixed-order calculation, although the agreement with NNLOJet is considerably better for and than for and , with the latter showing the largest differences of up to 20%.
@@ IV.2 Matching-scheme variations
Having observed very good agreement between different parton-shower prediction and the fixed-order calculation, we now investigate ambiguities in the matching scheme. Since different matched parton-shower predictions generally agreed very well, we focus on Powheg Box + Pythia, where we expect the largest differences arising from ambiguities in the multiplicative POWHEG approach and the interfacing to an external shower generator. Specifically, we study the influence of the following Powheg Box event-generation parameters:
@@ IV.4 Comparison of inclusive and exclusive setups
Finally, we contrast inclusive and exclusive predictions in the fiducial phase space. Specifically, we consider the Sherpa MEPS@NLO results with the MiNNLO + Pythia and MiNNLO + Herwig predictions as prototypes of inclusive results and compare them to the NLO predictions obtained with Sherpa MC@NLO, Powheg Box + Pythia, and Powheg Box + Herwig. Caution is advised in interpreting this comparison, as the MiNNLO predictions only achieve formal LO accuracy for configurations (and the logarithmic accuracy in case of MiNNLO + Herwig might not even be leading logarithmic), while all other setups, including the Sherpa MEPS@NLO prediction, are NLO precise with leading logarithmic accuracy on the parton shower side.
Predictions for the , , and are shown in Figs. 6-9. There is some difference between modeling higher jet multiplicities at leading order (Sherpa MEPS@NLO) and treating them as part of the inclusive calculation (Sherpa MC@NLO, Powheg Box + Pythia, Powheg Box + Herwig), in particular for observables that are sensitive to additional hard jets, such as . For variables such as , and , differences are smaller.
Perhaps the most striking feature is observed in the comparison of , shown in Fig. 9. There is a marked shape difference between MiNNLO and MEPS@NLO in the region around , a region important for studies of CP-violation, for both VBF Higgs and ggF production. This difference is amplified at the parton level. Moreover, there is a dramatic shape difference between MiNNLO + Herwig and MiNNLO + Pythia, with MiNNLO + Herwig indicating a similar behavior around as the NLO matched predictions. As a consequence of these observations, we conclude that the MiNNLO simulation has significant deficiencies when it comes to predicting the azimuthal correlations between the leading jets. This adds to the problems of describing the leading jet transverse momentum spectra observed in Fig. 2. On the other hand, we observe that all calculations that are NLO precise for the final state are consistent in their predictions of the spectrum. This includes in particular the MC@NLO predictions from Sherpa, the Powheg Box + Pythia and Powheg Box + Herwig predictions, and the inclusive MEPS@NLO predictions from Sherpa. We have verified that the MEPS@NLO results also agree with LO merged results from Sherpa, and that the latter are stable with respect to the number of jets included in the merging, as long as a third jet is included. Our recommendation is therefore to rely on any one of those predictions by default, and to gauge the uncertainty of the result by varying among the NLO accurate setups presented in this study.
@@ V The transverse momentum of the Higgs-plus-dijet system
Figure 10 (left) shows the results for NLO matched calculations. The agreement between most of the predictions is at the level of 10%, which is well within the range expected from renormalization scale variations for this complex final state. The largest difference is observed between the “simple shower” result from Powheg Box + Pythia and the prediction from Powheg Box + Herwig. In this case, the deviation is about 20%, which can still be considered a good agreement, given the complexity of the Born process.
@@ VI Cross section in cut based analyses
In Fig. 11, we show the cross sections for a number of predictions which are NLO accurate in the two jet phase space, compared to the prediction from MiNNLO. Bin 0 shows the inclusive two-jet prediction, satisfying the rapidity and transverse-momentum cuts.333The Sherpa MC@NLO prediction in Fig. 11 left does not exactly reproduce the inclusive rate of the parton-level result in Fig. 11 right (Bin 0 of the histogram). This is due to the generation cuts and the very loose jet requirements, which result in a larger effect from jet production through multiple parton scatterings than in Sherpa MEPS@NLO and Powheg Box+Pythia. A more inclusive simulation (such as Sherpa MEPS@NLO) is needed if one wishes to resolve the final state in this region of the phase space. The left panel shows the results at the hadron level with MPI, and the right panel for different NLO+PS setups at the parton level. All predictions agree well, as expected given the normalization conditions described earlier. In bins 1-6, minimum cuts are placed on the dijet mass of 60, 120, 350, 700, 1000 and 1500 GeV. There is reasonable agreement with MiNNLO (on the order of 10%), but MiNNLO predicts higher invariant mass final states than the other simulations. In bins 7-11, dijet rapidity separation cuts of 1, 2.0, 2.5, 3.0, 3.5 are applied. There is good agreement among the predictions, although Sherpa MC@NLO predicts smaller rapidity separations than the other simulations.

### v2

@@ III.1 Observable definitions
Additional data are collected on https://doi.org/10.5281/zenodo.17015614, together with the Rivet Buckley et al. (2013); Bierlich et al. (2020, 2024) analysis routine. A selection of further figures not contained in the main text below is given in Appendix A.
@@
The jets are clustered using the FastJet implementation of the anti- algorithm Cacciari et al. (2008), and must satisfy the transverse momentum cut .222We have verified that the conclusions of our study are unchanged when working with more exclusive phase-space cuts, in particular a Higgs boson rapidity cut of and a jet rapidity cut of . We primarily consider jets with a radius of 0.4, but also provide results for radii of 0.3, 0.6 and 0.8, to quantify the impact on selected differential observables.333Our study of the VBF Higgs boson production process in Ref. Buckley et al. (2021) examined jet -values from 0.3 to 1.0, and indicated a strong -dependence due to the kinematic requirements on the two jets. The R-dependence increased upon the imposition of the VBF phase space cuts, so it is important to check if the same effects occur for the background. Note that the two leading jets in VBF Higgs boson production are primarily quark jets, while in ggF they are primarily gluon jets. This may led to some differences in the -dependence. Jets are ranked by the magnitude of their transverse momentum. As a diagnostic tool, we sometimes also investigate observables in the Higgs+1 jet selection, which is defined as follows: Jets are clustered using the FastJet implementation of the anti- algorithm Cacciari et al. (2012), and must satisfy .
@@
To distinguish between normalization and shape differences of the predictions, all results will be normalized to the inclusive cross section for the final state, defined at the particle level, see Appendix A. This means in particular that a normalization difference can occur through the different behavior of the parton shower simulation, for example if the parton shower shifts a jet above/below the cut by means of additional radiation and/or recoil effects. The cross section is inferred from a relatively inclusive distribution, which we choose to be the Higgs boson rapidity distribution in the two jet final state.
III.2 Results of initial comparison
@@
We are now ready to compare results from the MiNNLO simulation as used by ATLAS and CMS to those from the tools introduced in Sec. II. The MiNNLO calculation is matched to the default Pythia parton shower (“simple shower”), both with and without the “dipole-recoil” option, and to the Herwig angular-ordered parton shower. The results are compared to a MEPS@NLO simulation from Sherpa, including the , , and final states at NLO precision, and up to at LO accuracy. This sample will be dubbed MEPS@NLO in the following. All predictions are at parton level, without multi-parton interactions and non-perturbative corrections. We also use a fixed-order NNLO calculation from NNLOJet as a reference.
@@
We find approximate agreement of the rate in the Higgs rapidity distribution of the selection (cf. Fig. 14) between NNLOJet and MiNNLO + Pythia, whereas the Sherpa MEPS@NLO sample requires a scale factor of , and the MiNNLO + Herwig sample a scale factor of . In the case of the Sherpa prediction, the scale factor is consistent with the NNLO to NLO ratio of the inclusive Higgs boson production cross section Anastasiou et al. (2016), which is needed to bring the inclusive jet rates in the MEPS@NLO result into approximate agreement with the prediction from NNLOJet. We will make use of these scale factors in most of the following comparisons.
@@
In Fig. 2 we show the leading and sub-leading jet transverse momentum spectra in the inclusive selection. The reference prediction stems from NNLOjet, which provides NLO precision for this observable. The MiNNLO simulation models it at LO precision. The MEPS@NLO simulation predicts it at NLO precision. To gauge the difference between LO and NLO accurate simulations, we also show a leading-order multi-jet merged result, labeled MEPS@LO. The MEPS@NLO, MEPS@LO and MiNNLO results all include Sudakov suppression effects, and both MEPS@NLO and MEPS@LO include higher-order tree-level corrections up to jets. Figure 2 shows that the shape of the scaled MEPS@NLO and MiNNLO +Pythia results agree with the fixed-order calculation of the leading jet transverse momentum within over the entire fiducial phase space, and in the medium hard region in particular. Deviations between MEPS@NLO and MEPS@LO are of similar size at transverse momenta below , but they can reach about 10% in the hard region . The shape of the MiNNLO +Pythia (dip) and MiNNLO +Herwig predictions differs significantly from the fixed-order result and the other two predictions. The situation is different for the sub-leading jet transverse momentum, where the MEPS@NLO and MiNNLO +Herwig prediction agree with the fixed-order result from NNLOjet within five percent up to GeV, and within about 10% up to 400 GeV. Here, the MiNNLO +Pythia and MiNNLO +Pythia (dip) predictions show deviations up to 20% already in the region below GeV. We will discuss similar features in more detail in Sec. IV.4, and in particular when comparing the distributions in Fig. 11.
@@
Figure 3 shows the comparison of the NNLOJet reference to the MEPS@NLO and MiNNLO predictions, which will be referred to as inclusive event-generation setups in the following. The middle panels of each sub-figure display the ratio to the inclusive fixed-order calculation from NNLOJet, while the bottom panels compare the event-generator predictions, normalized to the result from MiNNLO + Pythia with default recoil. The rapidity separation and the invariant mass distribution of the two leading jets, shown in Figure 3, agree well between the Sherpa, Pythia, and Herwig setups. All results also agree well with NNLOJet. The distributions of the azimuthal separation generally show good agreement between the different predictions in the region of small and intermediate separations, where they lie within a 10% envelope of each other. At large jet separations, , the MiNNLO + Pythia prediction is larger than all other predictions by about 30%, including the MiNNLO + Pythia results using the dipole-recoil option. We will discuss this in more detail in Fig. 11. The Higgs transverse-momentum distribution shows the largest shape differences between the different simulations. At low transverse momentum, , both MiNNLO + Pythia results and MiNNLO + Herwig exceed the NNLOJet result by about 20%, with close agreement between the MiNNLO predictions obtained with Herwig and with the Pythia dipole-recoil shower. The Sherpa MEPS@NLO prediction is about 20% lower than the NNLOJet result. Starting from , both MiNNLO + Pythia predictions agree very well with NNLOJet, whereas the Sherpa MEPS@NLO and MiNNLO + Herwig predictions create an envelope around the fixed-order result, with Sherpa MEPS@NLO being larger by about 10% and MiNNLO + Herwig being smaller by about the same amount. As can be inferred from the middle panes in all four figures, the event-generator predictions broadly lie within the seven-point scale variation uncertainty band of the fixed-order calculation, with some exceptions towards larger values of the observables.
Considering the observations in Figs. 2 and 3, one may be tempted to conclude that the modeling of the irreducible background to VBF Higgs production suffers from large theory uncertainties. This is in fact not the case. The event-generator predictions compared so far have rather different formal accuracy in the fiducial phase space, and therefore provide a different level of theoretical control over the observables of interest. In particular, the MiNNLO and MEPS@LO results are formally leading order accurate for these observables. However, the MiNNLO results differ between Herwig and Pythia for all observables in a way that is not suggested by the difference between the MEPS@LO and the MEPS@NLO results. We are therefore led to the conclusion that the difference in the simulations is not simply due to a difference in fixed-order higher-order corrections. We will focus on a practical solution to mitigate such problems with modeling the final state, which also relies on the simulations with the highest parametric precision for the jet final state.
@@ IV Assessment of theory uncertainties
In the following, we will use a consistent baseline for the hard, perturbative calculation – at NLO precision – and assess the influence of different theory variations on the size of the theoretical uncertainties in the fiducial phase space. Throughout the discussion, we will focus on the same four observables as in Fig. 3, with additional results available at https://doi.org/10.5281/zenodo.17015614. In Sec. IV.1, we compare parton-shower variations in NLO-matched distributions at the parton level. Sec. IV.2 focuses on matching variations in Powheg Box, while the influence of dedicated Pythia tunes and different hadronization models in Sherpa is discussed in Sec. IV.3. A study of the impact of the Higgs effective theory approximation can be found in App. C and a study of the jet-radius dependence can be found in App. D.
@@ IV.1 Parton-shower variations
As before, we scale all distributions to the NNLOJet rate based on the Higgs rapidity spectrum, yielding scale factors of in the case of Sherpa, for both Pythia showers, and for Herwig. The resulting parton-shower variations are shown in Fig. 5, including the NNLOJet NNLO prediction as a baseline. Besides the main distributions, we also show two ratio panels, with the first (middle panel) comparing the matched predictions to the NNLOJet fixed-order reference, and the second (bottom panel) showing the difference between the matched parton-shower predictions, normalized to the Powheg Box + Pythia result. All predictions in Figs. 4 and 5 acquire formal NLO accuracy in configurations.
Figure 4 shows similar agreement between the NLO matched results as between MEPS@NLO and MiNNLO in Fig. 2. As can be seen from Fig. 5, we find percent-level agreement between the four matched parton-shower algorithms for the , , and distributions. For the distribution, both Pythia showers agree very well with Sherpa, while the Herwig angular-ordered shower differs from these at large transverse momentum, . Except for the Higgs transverse-momentum distribution, the matched parton-shower results are within the scale uncertainties of the fixed-order calculation, although the agreement with NNLOJet is considerably better for , and than for , with the latter showing the largest differences of up to 20%.
@@ IV.2 Matching-scheme variations
We now turn to investigate ambiguities in the matching scheme. Since different matched parton-shower predictions generally agreed very well, we focus on Powheg Box + Pythia, where we expect the largest differences arising from ambiguities in the multiplicative POWHEG approach and the interfacing to an external shower generator. Specifically, we study the influence of the following Powheg Box event-generation parameters:
@@ IV.4 Comparison of inclusive and exclusive setups
Finally, we contrast inclusive and exclusive predictions in the fiducial phase space. Specifically, we consider the Sherpa MEPS@NLO results with the MiNNLO + Pythia and MiNNLO + Herwig predictions as prototypes of inclusive results and compare them to the NLO predictions obtained with Sherpa MC@NLO, Powheg Box + Pythia, and Powheg Box + Herwig. Caution is advised in interpreting this comparison, as the MiNNLO predictions only achieve formal LO accuracy for configurations (and have been designed for matching with transverse-momentum ordered showers), while all other setups, including the Sherpa MEPS@NLO prediction, are NLO precise with leading logarithmic accuracy on the parton shower side.
Predictions for , , , and are shown in Figs. 8-11. There is some difference between modeling higher jet multiplicities at leading order (Sherpa MEPS@NLO) and treating them as part of the inclusive calculation (Sherpa MC@NLO, Powheg Box + Pythia, Powheg Box + Herwig), in particular for observables that are sensitive to additional hard jets, such as . For variables such as , and , differences are smaller.
Perhaps the most striking feature is observed in the comparison of , shown in Fig. 11. There is a marked shape difference between MiNNLO and MEPS@NLO in the region around , a region important for studies of CP-violation, for both VBF Higgs and ggF production. This difference is amplified at the parton level. Moreover, there is a dramatic shape difference between MiNNLO + Herwig and MiNNLO + Pythia, with MiNNLO + Herwig indicating a similar behavior around as the NLO matched predictions. As a consequence of these observations, we conclude that the MiNNLO simulation has certain deficiencies when it comes to predicting the azimuthal correlations between the leading jets. This adds to the problems of describing the leading jet transverse momentum spectra observed in Fig. 2. On the other hand, we observe that all calculations that are NLO precise for the final state are consistent in their predictions of the spectrum. This includes in particular the MC@NLO predictions from Sherpa, the Powheg Box + Pythia and Powheg Box + Herwig predictions, and the inclusive MEPS@NLO predictions from Sherpa. We have verified that the MEPS@NLO results also agree with LO merged results from Sherpa, and that the latter are stable with respect to the number of jets included in the merging, as long as a third jet is included.
@@ V The transverse momentum of the Higgs-plus-dijet system
Figure 12 shows the results for NLO matched calculations. The agreement between most of the predictions is at the level of 10%, which is well within the range expected from renormalization scale variations for this complex final state. The largest difference is observed between the “simple shower” result from Powheg Box + Pythia and the prediction from Powheg Box + Herwig. In this case, the deviation is about 20%, which can still be considered a good agreement, given the complexity of the Born process.
@@ VI Cross section in cut based analyses
In Fig. 13, we show the cross sections for a number of predictions which are NLO accurate in the two jet phase space, compared to the prediction from MiNNLO. Bin 0 shows the inclusive two-jet prediction, satisfying the rapidity and transverse-momentum cuts.444The Sherpa MC@NLO prediction in Fig. 13 left does not exactly reproduce the inclusive rate of the parton-level result in Fig. 13 right (Bin 0 of the histogram). This is due to the generation cuts and the very loose jet requirements, which result in a larger effect from jet production through multiple parton scatterings than in Sherpa MEPS@NLO and Powheg Box+Pythia. A more inclusive simulation (such as Sherpa MEPS@NLO) is needed if one wishes to resolve the final state in this region of the phase space. The left panel shows the results at the hadron level with MPI, and the right panel for different NLO+PS setups at the parton level. All predictions agree well, as expected given the normalization conditions described earlier. In bins 1-6, minimum cuts are placed on the dijet mass of 60, 120, 350, 700, 1000 and 1500 GeV. There is reasonable agreement with MiNNLO (on the order of 10%), but MiNNLO predicts higher invariant mass final states than the other simulations. In bins 7-11, dijet rapidity separation cuts of 1, 2.0, 2.5, 3.0, 3.5 are applied. There is good agreement among the predictions, although Sherpa MC@NLO predicts smaller rapidity separations than the other simulations.